Cane Company manufactures two products called Alpha and Beta that sell for $210 and $172, respectively. Each product uses only one type of raw material that costs $8 per pound. The company has the capacity to annually produce 128,000 units of each product. Its unit costs for each product at this level of activity are given below: Alpha Beta Direct materials $ 40 $ 24 Direct labor 38 34 Variable manufacturing overhead 25 23 Traceable fixed manufacturing overhead 33 36 Variable selling expenses 30 26 Common fixed expenses 33 28 Total cost per unit $ 199 $ 171 The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are deemed unavoidable and have been allocated to products based on sales dollars. Required: Assume that Cane’s customers would buy a maximum of 98,000 units of Alpha and 78,000 units of Beta. Also assume that the company’s raw material available for production is limited to 248,000 pounds. What is the maximum contribution margin Cane Company can earn given the limited quantity of raw materials? Total Contribution Margin $
Answer #1

Contribution margin per pound

Alpha Beta
Selling price 210 172
Material 40 24
Direct labor 38 34
variable manufacturing overhead 25 23
Variable selling expense 30 26
Contribution margin per unit 77 65
Pound per unit 5 3
Contribution margin per pound 15.4 21.67

Pound used in beta = 78000*3 = 234000

Pound used in alpha (248000-234000) = 14000/5 = 2800 Units

Maximum contribution margin = (78000*65+2800*77) = $5285600